What is the electric flux passing through a Gaussian surface that contains 2 protons and 1 electron?


1
Expert's answer
2022-02-07T09:55:31-0500

The Gauss's law says


"\phi=\oint {\bf E}d{\bf A}=\frac{q_{\rm net}}{\epsilon_0}"

Hence, the electric flux passing through a Gaussian surface that contains 2 protons and 1 electron


"\phi=\frac{2e-e}{\epsilon_0}=\frac{e}{\epsilon_0}\\
=\frac{1.6*10^{-19 }}{8.85*10^{-12}}=1.8*10^{-8}\:\rm N\cdot m^2/C"